please help me solve: If the length of a rectangle is 2 more than its width, and its area is 35 square feet, what are the measures of the length and width?
Answer by addingup(3677)

begin with the formula for the area
L*W = area
And the problem says:
L = W+2
So now we have:
(W+2)*W = 35
W^2+2W-35 = 0
(W-5)(W+7)= 0
Split into 2 equations:
W-5 = 0
or
W+7 = 0
W = 5
or
W= -7 We know we're looking for a positive number, so toss this one out. Your answer is:
W = 5
`````````````````````````````````````````````````
Check:
(5+2)*5 = 35
7*5 = 35 Correct
John